OpenStudy (anonymous):

Find the minimum surface area for a 1 liter cylinder Use x for the radius

OpenStudy (briannabug):

Surface Area = two times the radius squared times pi plus two times the radius times the height times pi. Pi ( ) is approximately equal to 3.14 Or 22 7 Units of measure are square.

OpenStudy (briannabug):

01FormulaReferenceSheet10.pdf

OpenStudy (anonymous):

Umm yes I already have all that stuff but I need a definitive way to find the values of the radius and height to get the least possible surface area I already know all the formulas and stuff

OpenStudy (anonymous):

Some sites I've visited suggested finding the derivative but I'm only in Pre-calc and derivatives are in Calculus

OpenStudy (briannabug):

Okay do you have a pic of the cylinder or no? Just wondering if not its okay.

OpenStudy (anonymous):

No it provides no picture But it's really simple, it's a cylinder with a radius of x, that is all the information provided

OpenStudy (briannabug):

Okay, as long as you have your radius and height, you can go ahead and plug those values into your formula to find your answer for surface area. The formula for surface area gives you the total area of all the surfaces together.

OpenStudy (anonymous):

But you see I'm not given a radius or height, What I am given is the volume, I need to find the radius and height through a method that is preferably not guess and check

OpenStudy (briannabug):

Okay just give me a minute and I will show you how. Is that okay?

OpenStudy (briannabug):

First of all, what is the volume?

OpenStudy (briannabug):

???

OpenStudy (briannabug):

I need the volume to continue to help you.

OpenStudy (anonymous):

It's 1 liter, I said it in the problem

OpenStudy (briannabug):

Okay

OpenStudy (briannabug):

Is it multiple choice of type in?

OpenStudy (briannabug):

Or neither???

OpenStudy (anonymous):

There is no multiple choice THis is an open ended question, There are ways to solve this but I prefer not to use guess and check or graphing

jhonyy9 (jhonyy9):

@Brill how you get for a cylinder volume equal 1 liter what dimensions need having the cilinder for you get volume equal 1 liter - do you know it ?

OpenStudy (anonymous):

Well you can covert it into cubic centimeters, where 1000 cubic centimeters equals 1 liter

OpenStudy (anonymous):

Don't know if that's what you wanted

OpenStudy (briannabug):

@Brill @jhonyy9 Volume of a cylinder is in other words the capacity of the shape, with the assumption that its walls are extremely thin. It is often expressed in cubic meters (m³) or cubic centimeters (cm³) (1 cm³ = 1 ml). In everyday life it is easier to find the use of liter unit, which is equivalent to 1,000 cm³. A standard sized can used by most of the fizzy drinks manufacturers is in approximation a cylinder with capacity of 330 ml = ⅓ of a liter. Now when we know what the volume of a cylinder is it would be good to know how to calculate it without the necessity of filling it up with liquid. The calculation turns out to be quite straightforward and requires calculating the area of a circle that forms the base of a cylinder and multiplying the result by the height of that cylinder

OpenStudy (briannabug):

Okay, I don't know if any of this helped but if it did just let me know.

OpenStudy (anonymous):

Yes I know @briannabug I already learned that in geometry But I'm not sure how to calculate the least radius and height

OpenStudy (anonymous):

Well not the least radius and height but the least surface area

jhonyy9 (jhonyy9):

@Brill so using the above wrote that 1000 cm^3 mean for volume 1 liter you can us it for calculi the dimensions of cylinder for you get like result of volume 1000 cm^3 what mean 1 liter - do you understand it now ?

jhonyy9 (jhonyy9):

you know that the volume of a cylinder you get from multiplie area of base time height - so but you need get it for least radius area of base what is a circle and area of circle is pi*r^2 sorry are you here ?

OpenStudy (anonymous):

Yeah I am sorry Go on please

OpenStudy (anonymous):

Umm sorry what do you mean by "you can us it for calculi the dimensions of cylinder for you get like result of volume 1000 cm^3 what mean 1 liter"

jhonyy9 (jhonyy9):

so 1 liter = 1000 cm^3= 3,14 *r^2 *10 let height of cylinder =10 cm r^2 = 1000/31,4 r^2 = 31,84 r = sqrt(31,84 = 5,64 cm radius = 5,64 cm

OpenStudy (anonymous):

Ok but why do we make the height equal 10? Is this not just the same as guess and check?

jhonyy9 (jhonyy9):

so for radius 5,64 cm and height of cylinder =10 cm you need to calculi now the total surface area of cylinder - hope you can do it - yes ?

jhonyy9 (jhonyy9):

this 10 cm for height is just one example for total surface area - after you calculi this so than you can check it how you can making it smaller than you make smaller what dimensions - ok ?

OpenStudy (anonymous):

Yeah that's 554.237263 cm squared

OpenStudy (phi):

****Find the minimum surface area for a 1 liter cylinder Use x for the radius ****** the volume of the cylinder is area of the base times its height: \[ V= \pi x^2h=1 \] from which we get \( h= \frac{1}{\pi x^2} \) the total surface area is the area of the top and bottom i.e. 2 pi x^2 plus its side: \[ A = 2 \pi x^2 + 2 \pi x h \] or, replacing h: \[ A= 2 \pi x^2 + \frac{2}{x} \] that is the surface area of a cylinder as a function of its radius (but with fixed volume=1) Here is a graph of part of the curve

OpenStudy (phi):

To find the min value on that curve, take the derivative of A with respect to x and set it equal to zero. \[ \frac{dA}{dx} = 4 \pi x -2 x^{-2}= 0 \] solve for x, and use that value to find A

OpenStudy (anonymous):

See the problem with using derivatives is that I haven't gotten to them yet in classes so I have no idea what they are, Is there another way to do this?

OpenStudy (phi):

Yes, another way is to look at the graph and find the min value, which you can estimate

OpenStudy (anonymous):

ok thank you very much!